OpenStudy (anonymous):

Hello! Please will someone help me with the following question, I am going wrong somewhere. Find the co-ordinates of the points of intersection of the line y=2x and the curve y=x^2+6x-5.

OpenStudy (anonymous):

Here is my working out so far. I didn't carry on because I know theres already something wrong.

OpenStudy (anonymous):

If you're completing the square, the number you add is \[\left( \frac{ b }{ 2 } \right)^2\]. Your b is 4, so you should add (4/2)² = 4 to both sides. Your third line should be \[x^2+4x+4=9\]

OpenStudy (anonymous):

Oh I see! Thank you!

OpenStudy (anonymous):

you're welcome